OBSERVATION OF THE ROTATIONAL SPECTRA OF ^{4}HeH^{+}, ^{4}HeD^{+}, ^{3}H eH^{+}, and ^{3}HeD^{+}

\underline{KOJIRO TAKAGI}, FUSAKAZU MATSUSHIMA, Department of Physics, Toyama University, Toyama 930, Japan; TAKESHI OKA, Department of Chemistry, Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637.

Low J rotational transitions of ^{4}HeH^{+}, ^{4}HeD^{+}, ^{3}HeH^{+}, and ^{3}HeD^{+} have been observed in the 2-5 THz frequency region with a high-precision far-infrared spectrometer using a tunable radiation source. Observed frequencies have been incorporated with previous infrared measurements to determine Dunham coefficients Y_{kl} and isotopically independent parameters U_{kl}, Δ_{kl}^{He}, and Δ_{kl}^{H}. In particular, constants with k=0 and l=1 and 2 have been determined with unprecedented accuracy and will provide important information for breakdown of the Born-Oppenheimer approximation. The lowest J=1←0 transition of ^{4}HeH^{+} observed at 2010.1839(2) GHz will be an important future probe for detecting the species in space.